RustedKatana:

Solve 2x2 − 9x + 7 = 0 using an appropriate method. Show the steps of your work and explain why you chose the method used.

RustedKatana:

@sunflower0kiddo or @VIBE

Sunflower0Kiddo:

I gocchu

Sunflower0Kiddo:

Let's solve your equation step-by-step, 2x2−9x+7=0

RustedKatana:

Ok.

RustedKatana:

What do I need to do first?

Sunflower0Kiddo:

Step 1: Factor left side of equation, (2x−7)(x−1)=0

RustedKatana:

Ok, how do I work that out?

Sunflower0Kiddo:

Step 2: Set factors equal to 0. 2x−7=0 or x−1=0 x= 7/2 (fraction) or x=1

Sunflower0Kiddo:

Just copy the numbers ^^

RustedKatana:

Ok. What next?

snowflake0531:

@sunflower0kiddo wrote:
Step 1: Factor left side of equation, (2x−7)(x−1)=0
Show how you got that `Show the steps of your work and explain why you chose the method used.`

RustedKatana:

How do I do that though?

Joe348:

wait is 2x2 actually 2x^2 ?

RustedKatana:

yes

Joe348:

okay

Sunflower0Kiddo:

I used the website i told him/her to use :P

RustedKatana:

I did but now it says I have to upgrade to use it.

snowflake0531:

Which website are you guys talking about? mathway? wolframalpha?

RustedKatana:

mathpapa.com

Sunflower0Kiddo:

mhm

Joe348:

mathway is not a good webstie for math ngl

RustedKatana:

yeah ik

Sunflower0Kiddo:

true

snowflake0531:

Well, you sure you can use calculators, but you have to still learn how to do it by yourself .-.

Sunflower0Kiddo:

i suck at intigers-

Joe348:

Use the quadratic formula in this equation

snowflake0531:

o-O It would be easier to just do regular factoring in this one

RustedKatana:

Ok, Thank you! :-)

Sunflower0Kiddo:

:>

Sunflower0Kiddo:

@joe348 Hey, can i snag you for help in Dm's?

Joe348:

Isnt there actually 2 X snow?

snowflake0531:

yes there is

snowflake0531:

what are you getting at-?

Joe348:

No I was just making sure there was

Joe348:

Because I kinda found it

snowflake0531:

found what

Sunflower0Kiddo:

@snowflake0531 wrote:
found what
i think he means the answer

snowflake0531:

well, @sunflower0kiddo already gave the answer earlier lol

Joe348:

Yeah but I dont wanna tell you straight up so here how to solve it idk if this help but yeah

Sunflower0Kiddo:

@snowflake0531 Oh my god- Im fricking stupid

Sunflower0Kiddo:

brb

snowflake0531:

-?

Joe348:

did my answer even help well o the way to the answer? on*